# Optimized Multi-Token Joint Decoding with Auxiliary Model for LLM Inference

### v2

@@ Appendix A Proof
Appendix B Pseudocode of MJSD
@@ Appendix C Additional Experiments
C.1 Resutlts with OPT-30B and Llama-2-70B
Here we report the performances of different methods for OPT (350M and 30B) and Llama-2-Chat (7B and 70B). Table 9 shows the average performances across all datasets. MTAD always achieves the lowest perplexity and the best efficiency.
@@ C.1 Resutlts with OPT-30B and Llama-2-70B
C.2 Ablation study of top-k and top-p sampling
Here we conduct an ablation study to show how the value of $k$ and $p$ in top-k and top-p warping affects our method. Table 10 shows the results. We can see that when changing the value of $k$ and $p$, MTAD consistently achieves significantly better performances.
Table 10: Ablation study of $k$ and $p$ in top-k and top-p sampling
K P Greedy Speculative MJSD
PPL Tokens/sec PPL Tokens/sec PPL Tokens/sec
20 0.9 3.74 22.6 3.64 36.8 2.06 63.0
20 0.8 3.06 22.7 3.10 38.5 1.93 58.8
10 0.9 3.03 22.7 3.22 38.5 1.95 62.5
10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
C.3 Additional Experiments on CNNDM and Spider
We also report the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset. Table 11 and Table 12 show the results. We can see that MTAD consistenly achieve better effectiveness as well as faster decoding speed.
@@ C.3 Additional Experiments on CNNDM and Spider
Speculative 11.5 28.5 11.5 27.8 16.3 25.6
MJSD 13.0 30.3 14.8 32.3 18.3 29.4
@@ Appendix D Energy Consumption Measurement
We use the command "nvidia-smi --query-gpu=power.draw --format=csv" to get GPU power every second, and sum them up as the total energy consumption. We use average energy consumption per token to measure energy efficiency. There is a recent study pointing out the measurement error using nvidia-smi (Yang et al., 2023). We follow the three principles proposed in (Yang et al., 2023) to minimize the error.
@@ Appendix E Configuration
The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving (Schuster et al., 2022). We set the maximum running time to be an hour for each baseline. We use average tokens/second to measure the inference speed and use perplexity (exponentiated average negative log-likelihood) based on the probability of the large model to measure the output quality. Because different methods might finish different numbers of inputs, we only calculate the perplexity of the first $M$ inputs, where $M$ is the number of inputs finished by greedy decoding. We use average energy consumption per token to measure energy efficiency. The details of energy measurement are illustrated in the Appendix.
@@ Appendix F Hyper-parameter Details
In the experiments, we follow the default settings in (Bear, 2024) to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warpping operations in a public speculative decoding implementation.
  1. 1.
    Keep the probabilities of top 20 tokens unchanged, and set the probabilities of other tokens to 0, then normalize the distribution.
  2. 2.
    Sort the tokens based on their distributions descendingly. Keep the first $K$ tokens such that their cumulative probabilities is larger than 0.9, while set the probabilities of other tokens to 0.
For different methods, we choose their hyper-parameters by using a small validation set. We select the set of hyper-parameters that make the corresponding method have best output perplexity. Table 13 shows the hyper-parameters used in the experiments.
Table 13: Hyper-parameters of different methods for different models and datasets. L: Llama, O:OPT, CP: ChatGPT-Prompts, CA: ChatAlpaca, CD: CNNDaily.
L,CP O,CP L,CA O,CA L,CD O,CD L,SP O,SP L,MT O,MT
speculative step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
Spectr&SpecInfer step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
num seq $m$ 4 4 2 2 4 2 2 2 2 2
BiLD step len $\gamma$ 10 10 10 10 10 10 10 10 10 10
fallback thres $\tau_{1}$ 0.9 0.9 0.9 0.3 0.9 0.3 0.9 0.9 0.9 0.9
rollback thres $\tau_{2}$ 2 2 1 2 3 2 1 1 1 1
MTAD step len $\gamma$ 4 4 4 4 4 4 4 4 4 4
num beams 8 8 8 8 8 8 8 8 8 8
acc/rej thres $\tau$ 0.1 0.1 0.1 0.1 0.1 0.1 0.9 0.1 0.9 0.1

### v3

@@ Appendix A Proof
Appendix B Pseudocode of MTAD
@@ Appendix C Additional Experiments
[…]
Llama-2 OPT
SpD 2.02$\pm$0.05 2.60$\pm$0.06
BiLD 1.83$\pm$0.10 2.68$\pm$0.36
Spectr 2.73$\pm$0.43 3.45$\pm$0.42
SpecInfer 2.74$\pm$0.46 3.45$\pm$0.40
MTAD 3.17$\pm$0.43 4.30$\pm$0.03
SpD MTAD
CD Rouge-L 0.114 0.118
SP EA 11.5 13.0
MT Humanities 2.95 3.15
Extraction 1.80 2.50
Roleplay 3.10 3.80
Math 1.10 1.00
Coding 1.25 1.10
Reasoning 3.80 3.15
STEM 2.85 3.10
Writing 3.80 3.65
Average 2.58 2.68
C.1 Additional Datasets and Model Family
Here we report the additional experiment results with three more datasets: (1) ChatGPT-Prompt (Rashad, 2023), (2) ChatAlpaca (Bian et al., 2023), (3) CNN Dailymail (See et al., 2017). We use two public LLM families in our experiments: OPT (Zhang et al., 2022) and Llama-2 (Touvron et al., 2023). We set the large model to be OPT-13B and Llama-2-13B as they are the largest models that can run on a single 40GB GPU, and utilize Llama-68M (Miao et al., 2023) and OPT-125M as the small models.
Table 7 shows the full evaluation results, and Table 9 displays the downstream performance. Table 8 depicts the average number of generated token per iteration for different algorithms. The experiment results demonstrate that MTAD achieves better efficiency, better perplexity, as well as better downstream performance.
Table 10: Number of Generated Tokens Per Iteration Across Thresholds
Threshold
0.1 0.3 0.5 0.7 0.9
Llama-3 4.72 4.43 4.12 3.93 3.64
Llama-3-Instruct 4.74 4.51 4.28 4.06 3.79
C.2 Ablation Studies on Acceptance Threshold
Next, we evaluate the effect of acceptance threshold $\tau$. Intuitively, when we increase $\tau$ from 0 to 1, the acceptance criterion becomes stricter, the efficiency drops while the output perplexity increases. Surprisingly, this expectation is only partially correct. As shown in Figure 3, the efficiency indeed drops when $\tau$ increases. However, the perplexity rises when $\tau$ is close to 1. When $\tau=1$, all the draft tokens are rejected, which makes MTAD equivalent to multinomial sampling. Similarly, when $\tau$ is close to 1, the advantage of multi-token joint decoding on effectiveness disappears, hence MTAD’s perplexity is similar to the perplexity of multinomial sampling. Moreover, when $\tau$ ranges from $0.1$ to $0.9$, the performance of MTAD is relatively stable, suggesting that MTAD is not sensitive to the acceptance threshold.
In addition, Table 10 exhibits how the number of accepted tokens change with thresholds. We can see that the number of accepted tokens grows smaller when the threshold increases, which explains the speed degradation.
C.3 Ablation Studies on Number of Beams
In Table 11, we investigate how the number of beams used in the beam decoding of the small model affects the inference performance. Increasing the number of beams boosts the quality of the draft tokens, which not only improves the output perplexity but also increases the average acceptance length and hence leads to better efficiency. But we can see that the increment slows down when the number of beams is too large. In addition, when the number of beams is too large, the inference cost of the small model will become too high.
Table 11: Effect of beam count on inference performance (ChatGPT-Prompts dataset). Arrows indicate optimal direction ($\uparrow$=higher better, $\downarrow$=lower better).
Number of Beams
2 4 6 8
Llama-2 Speed (tok/s) $\uparrow$ 55.9 59.9 60.2 61.3
Energy (J/tok) $\downarrow$ 2.43 2.25 2.22 2.20
Perplexity $\downarrow$ 2.44 2.12 2.14 2.10
OPT Speed (tok/s) $\uparrow$ 51.0 54.1 54.3 55.9
Energy (J/tok) $\downarrow$ 2.50 2.32 2.36 2.30
Perplexity $\downarrow$ 3.63 3.16 3.42 3.19
C.4 Ablation study of top-k and top-p sampling
Table 12 demonstrates how the value of k and p in top-k and top-p warping affects our method. We can see that by changing the value of $k$ and $p$, MTAD consistently performs significantly better.
Table 12: Ablation study of $k$ and $p$ in top-k and top-p sampling
K P Multinomial SpD MTAD
PPL Tokens/sec PPL Tokens/sec PPL Tokens/sec
20 0.9 3.74 22.6 3.64 36.8 2.06 63.0
20 0.8 3.06 22.7 3.10 38.5 1.93 58.8
10 0.9 3.03 22.7 3.22 38.5 1.95 62.5
10 0.8 2.56 22.7 2.53 40.0 1.80 62.5
C.5 Resutlts with OPT-30B and Llama-2-70B
Here we report the performances of different methods for OPT (350M and 30B) and Llama-2-Chat (7B and 70B). Table 13 shows the average performances across all datasets. MTAD always realizes the lowest perplexity and the best efficiency.
@@ C.5 Resutlts with OPT-30B and Llama-2-70B
C.6 Additional Experiments on CNNDM and Spider
Table 14 and Table 15 reveal the downstream effectiveness of our method on CNNDM and Spider when the small model and large model are fine-tuned on the dataset. We can see that MTAD is consistently more effective and has a faster decoding speed.
@@ C.6 Additional Experiments on CNNDM and Spider
SpD 11.5 28.5 11.5 27.8 16.3 25.6
MTAD 13.0 30.3 14.8 32.3 18.3 29.4
C.7 Visualization of perplexity and output quality
To further illustrate the relationship between perplexity and downstream performance, we present a scatter plot in Figure 4. The plot depicts the correlation between relative downstream scores (normalized by the score of multinomial sampling) and relative perplexity (normalized by the perplexity of multinomial sampling) across 7 decoding algorithms, 3 datasets, and 2 model configurations. The results confirm that lower perplexity generally correlates with higher output quality.
Figure 4: Relationship between relative perplexity (normalized by multinomial sampling’s perplexity) and relative performance score (normalized by multinomial sampling’s score).
C.8 Correlation between Energy and Speed
Figure 5 shows there is a correlation between speed and energy, whether we consider the entire table or focus on a specific dataset and model. For fairness, all methods for a given dataset and model were run on the same machine nodes. However, for a fixed method (e.g., Spectr), experiments on different datasets and models might be conducted on different nodes (all equipped with L40 GPUs). We did notice that the same configuration run on different machines may have varied energy consumption. This variation introduces some randomness, which could make the correlation appear less consistent across datasets and models.
(a) group by dataset and model
(b) group by methods
Figure 5: Correlation between speed and energy
@@ Appendix D Energy Consumption Measurement
To get GPU power every second, we run the command “nvidia-smi -query-gpu=power.draw -format=csv”. We add the results up to determine the total energy consumption. We use average energy consumption per token to measure energy efficiency. There is a recent study pointing out the measurement error using nvidia-smi (Yang et al., 2023). We follow the three principles proposed in (Yang et al., 2023) to minimize this error.
@@ Appendix E Configuration
The experiments are conducted on a machine with 1 Nvidia L40 GPU (48 GB), 4 CPUs, and 50 GB main memory, using a batch size of 1, which is common for online serving (Schuster et al., 2022). We set the maximum running time to be an hour for each baseline. We use average tokens/second to measure the inference speed and use average energy consumption per token to measure energy efficiency.
@@ Appendix F Hyper-parameter Details
In the experiments, we follow the settings in (Bear, 2024) to warp the sampling distribution $p$ and $q$ with the following steps, which are the default warping operations in a public speculative decoding implementation. Specifically, we first keep the probabilities of top 10 tokens unchanged, and set the probabilities of other tokens to 0, then normalize the distribution. Then we sort the tokens based on their distributions in descending order and keep the first $K$ tokens such that their cumulative probabilities is larger than 0.9, while set the probabilities of other tokens to 0.
For different methods, we choose their hyper-parameters based on a small validation set. We select the set of hyper-parameters that make the corresponding method have best output perplexity. For MTAD, we choose the beam width from $\{4,8\}$, the number of draft tokens from $\{3,4\}$, and the acceptance threshold from $\{0.1,0.3,0.5,0.7,0.9\}$.
@@ Appendix G License of Datasets and Models
  • MTBench: Apache-2.0 License
  • Spider: CC BY-SA 4.0 License
  • HumanEval: MIT License